OpenStudy (anonymous):

What is the equation of a line that passes through (-6, 2) and has a slope of −12 ?

OpenStudy (kidrah69):

\(\huge\color{orange}{Welcome~To~OpenStudy}\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://goo.gl/5pp1u0 For an visual tutorial look at this! Its amazing :o https://prezi.com/fs3hqdpcopic/an-unofficial-guide-to-openstudy/ Also consider reading the CoC and the T&C http://openstudy.com/code-of-conduct http://openstudy.com/terms-and-conditions

OpenStudy (kidrah69):

\[y-y1=m(x-x1) \] slope = -12 x1=-6 y1=2

OpenStudy (kidrah69):

Just plug everything in and solve :)

OpenStudy (anonymous):

y=x2−2 y=−x2−1 y=2x+1 y=−2x−1

OpenStudy (kidrah69):

Do what i told u to do above

OpenStudy (texaschic101):

in y = mx + b form the slope is in the m position so m = -12 your points (-6,2)....x = -6 and y = 2 now we already know the slope (m) and our points (x,y)....so basically we just need to find b, the y intercept. So we sub into the formula and find b y = mx + b 2 = -12(-6) + b 2 = 72 + b 2 - 72 = b -70 = b so your equation becomes : y = -12x - 70 questions ?

OpenStudy (anonymous):

thanks